Find the minor axis vertices of the ellipse.16x^2+4y^2+64x+24y+36=0
Answer by lwsshak3(11628)

Find the minor axis vertices of the ellipse.16x^2+4y^2+64x+24y+36=0
**
16x^2+4y^2+64x+24y+36=0
complete the square
16(x^2+4x+4)+4(y^2+6y+9)=-36+64+36
16(x+2)^2+4(y+3)^2=64
divide by 64
(x+2)^2/4+(y+3)^2/16=1
This is an equation of an ellipse with vertical major axis of the standard form:
(x-h)^2/b^2+(y-k)^2/a^2=1, a>b, with (h,k) being the (x,y) coordinates of the center.
..
For given ellipse:
center:(-2,-3)
b^2=4
b=√4=2
Vertices of minor axis=(-2±b,-3)=(-2±2,-3)=(-4,-3) and (0,-3)